Can someone help me with this one
Solve the system by graphing, addition and substitution methods.
5x – 4y = –5
x – 2y = –7

Answer by rapaljer(4671)

5x – 4y = –5
x – 2y = –7

Multiply the second equation by -2 to eliminate the y terms:
5x – 4y = –5
-2x +4y = 14

Add the equations:
3x=9
x=3

Substitute x=3 into the first equation:
5x – 4y = –5
5(3) -4y =-5
15-4y = -5
15-15-4y = -5-15
-4y = -20
y=5

Check by substituting both numbers into the second equation:
x – 2y = –7
3-2(5) = -7
It checks

Solution: (3,5)

R^2